OpenStudy (anonymous):

The length of a rectangle is 2 more than 3 times the width. If the perimeter is 100 meters, what is the width of the rectangle? 13 meters 12 meters 11 meters

OpenStudy (anonymous):

12 meters

OpenStudy (anonymous):

perimeter of rectangle=2(length+width) say, the width is x so, the length is 3x+2 so, perimeter=2(3x+2+x)=100 x=12

OpenStudy (anonymous):

@matricked , please provide explanation with ur answer. thanx :)

OpenStudy (anonymous):

http://openstudy.com/code-of-conduct